Circular No. 5670

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SWIFT-TUTTLE (1992t) Observations in mid-November show that ephemerides from the grav-

itational orbital elements on IAUC 5636 (by B. G. Marsden, utilizing the 1862 Oct. observations) and on MPC 21081 (by D. K. Yeomans, utilizing the 1862 July-Sept. observations) were both running off by as much as 15′′ in declination, the truth being roughly midway in between. As anticipated, gravitational solutions now completely fail to link the apparitions. The nongravitational solution below, from MPC 21235, satisfies 383 observations with a mean residual of 1′′.3 and removes much of the systematic effect, but the residual trend in 1862 Oct. is still in excess of 10′′ in both α and δ. The computed perihelion time in 1737 is ∼ 0.5 day too early.

Epoch = 1992 Dec. 4.0 TT

T = 1992 Dec. 12.3243 TT  ω = 153.^{o}0016
e = 0.963589  Ω = 139.4441 }2000.0
q = 0.958216 AU  i = 113.4266
a = 26.316841 AU n^{o} = 0.0073005 P = 135.005 years

Nongravitational parameters are A_{1} = −0.00079 ± 0.00002, A_{2} = −0.00266 ± 0.00007. The new solution also satisfies the following pre-recovery observations:

1992 UT  α_{2000}  δ_{2000}  m_{1} Observer
Jan. 3.07847 10^{h}00^{m}08.^{s}38 +42^{o}59′08′′.0 17.5 Haver
3.09184 10 00 07.84 +42 59 16.0  _{”}
7.08767 9 56 32.45 +43 56 49.2 18  Kohoutek

R. Haver (Cima Ekar). 0.4-m f/2.5 Schmidt. Measurers A. Boattini and M. Tombelli. Communicated by G. Forti. Very weak images; stellar appearance.

L. Kohoutek (Calar Alto). 0.8-m Schmidt. Close to plate limit; evidently

some interference from cirrus clouds.

Using computations by the undersigned and also by G. Waddington (Oxford University), G. Kronk (Troy, IL) has confirmed the likelihood of the –68 comet as a good candidate for P/Swift-Tuttle and has noted that the only other viable candidate is the comet of +188. He suggests that the observations restrict the perihelion dates to the ranges –68 July 27-28 and +188 July 5-Aug. 2 (with a most likely date of July 15).

1992 December 5 Brian G. Marsden